The following are likely effects created by a Minimum Wage Regulation, based on the model presented in class, EXCEPT:

Question options:

Increase in the number of workers looking for jobs

Decrease in the number of jobs available.

It will for sure increase wages received by the lowest skilled workers.

The following are possible characteristics related to different market structures, EXCEPT:

Question options:

Monopoly always implies the existence of positive economic profits

The key characteristic for Monopolistic Competition is product differentiation.

Collusion is a profitable but unstable strategy under oligopoly.

The following is the absolute value for the Mid-Point price elasticity of demand between two points on certain demand:

Point A : Po = 15 & Qo =10

while

Point B: P1 = 5 & Q1= 20

Question options:

l Ep l = 3/2

l Ep l = 2/3

l Ep l = 1/4
Answer #1

a)"C"

A minimum wage regulation will decrease the number of people employed and increase the number of people looking for job, it will decrease the wages received by the lowest skill worker.

b) "A"

Monopoly always doesn't imply a positive profit but sometimes they face loss too.

c) %change in demand = (20-10) / (20+10) /2 = 10 / 15= 1.5

%change in price (5-15) / (5+15) /2 = -10 / 10 = 1

answer is 1.5 / 1 or 3/2 "A".